# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 02 Mar 2026 v1

@@ 6.1 Local Fisher Geometry and a Frozen-Metric Regime
Suppose $F(\theta)$ is Lipschitz on a neighborhood $\mathcal{U}$ of $\theta^{\star}$: $\|F(\theta)-F(\theta^{\star})\|_{\operatorname{op}}\leq L_{F}\|\theta-\theta^{\star}\|_{2}$ for all $\theta\in\mathcal{U}$. If the SGD iterates satisfy $\|\theta_{t}-\theta^{\star}\|_{2}=o(1)$ almost surely, then for every $\epsilon>0$ there exists $t_{0}$ such that for $t\geq t_{0}$, $(1-\epsilon)F^{\star}\preceq F(\theta_{t})\preceq(1+\epsilon)F^{\star}$. In particular, the frozen-metric approximation $\|\cdot\|_{F^{\star}}$ and $\|\cdot\|_{F(\theta_{t})}$ are equivalent up to multiplicative $(1\pm\epsilon)$ for all sufficiently late iterates.
@@
The quantity $d_{\operatorname{eff}}(F^{\star})$ is exactly the stable rank of $F^{\star}$, a standard measure of intrinsic dimensionality that is robust to small perturbations (Tropp, 2015). Three properties make $d_{\operatorname{eff}}$ the appropriate dimensional parameter for Fisher-metric rates: (i) $d_{\operatorname{eff}}\leq d$ always, with equality if and only if all eigenvalues of $F^{\star}$ coincide (isotropic Fisher geometry); (ii) $d_{\operatorname{eff}}$ can be much smaller than $d$ when Fisher information is concentrated in a few directions—the extreme case $d_{\operatorname{eff}}\to 1$ occurs when a single direction dominates; (iii) $d_{\operatorname{eff}}$ replaces the ambient dimension $d$ in both the upper bound (Theorem 6.4) and the oracle complexity (Theorem 7.3). The distinction is operationally significant: classical Euclidean bounds scale with the ambient dimension $d$, whereas the Fisher-metric bounds scale with $d_{\operatorname{eff}}$, reflecting only the statistically distinguishable directions. In problems where information is concentrated (e.g., sparse or low-rank Fisher structure), $d_{\operatorname{eff}}\ll d$ and the Fisher-metric guarantee is correspondingly tighter.
@@ 6.2 Upper Bound
Remark 6.5 (OU viewpoint and the same $1/T$ scaling).
In the local regime, the diffusion limit yields a (time-rescaled) OU approximation whose diffusion coefficient is proportional to the frozen mini-batch covariance $F^{\star}/b$. With a diminishing schedule $\eta_{t}\asymp 1/t$, this perspective is consistent with the discrete-time bound: the mean-square error in the frozen Fisher norm decays on the order of $1/(Tb)$, with constants governed by $F^{\star}$ and the intrinsic curvature $\mu_{F}$.
@@ 6.3 Lower Bound
The oracle transcript $(\theta_{0},Y_{1},\dots,Y_{N})$ thus forms a martingale difference sequence with predictable quadratic variation bounded by $\mathcal{I}_{1}$.
Remark 6.7 (Beyond i.i.d. oracles).
The classical setting of i.i.d. sampling from $P_{\theta^{\star}}$ satisfies Assumption 6.6 with equality in (iii). However, the lower bound argument does not fundamentally rely on independence across oracle calls. What is required is a predictable quadratic-variation bound on the oracle transcript. This accommodates exchangeable or conditionally i.i.d. sampling schemes (e.g., de Finetti mixtures) provided the cumulative information growth is bounded by $N\mathcal{I}_{1}$. In this sense, the lower bound below is not a direct transposition of the classical i.i.d. van Trees proof but an oracle-information statement tailored to stochastic optimization.
Theorem 6.8 (Minimax lower bound in Fisher metric).
Suppose Assumption 6.6 holds and $F^{\star}:=F(\theta^{\star})\succ 0$. Let the parametric family $\{P_{\theta}:\theta\in\Theta_{\operatorname{loc}}\}$ satisfy the regularity conditions of the van Trees inequality on a neighborhood $\Theta_{\operatorname{loc}}$ of $\theta^{\star}$ (see Gill and Levit (1995); the connection to local asymptotic normality and intrinsic statistical distance is developed in Le Cam (1986); the key requirements are smoothness and integrability of the Fisher information on $\Theta_{\operatorname{loc}}$). Then
@@
Step 1 (predictable information accumulation). By Assumption 6.6(iii), the predictable quadratic variation of the oracle transcript satisfies
$\sum_{t=1}^{N}\mathbb{E}_{\theta^{\star}}\!\bigl[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}\bigr]\;\preceq\;N\,\mathcal{I}_{1}(\theta^{\star}).$
Thus the total predictable information in the oracle transcript is bounded by $N\mathcal{I}_{1}(\theta^{\star})$.
Step 2 (transcript measurability). Since the estimator $\hat{\theta}_{N}$ is $\mathcal{F}_{N}$-measurable, its Fisher information cannot exceed the predictable information accumulated in the oracle transcript.
Step 3 (van Trees inequality). Apply the van Trees inequality with a smooth prior $\pi$ supported on $\Theta_{\operatorname{loc}}$: $\mathbb{E}_{\pi}\mathbb{E}[(\hat{\theta}-\theta)(\hat{\theta}-\theta)^{\top}]\succeq(N\mathcal{I}_{1}+J_{\pi})^{-1}$. Multiply by $F^{\star}$ and take traces: $\mathbb{E}_{\pi}\mathbb{E}[\|\hat{\theta}-\theta\|_{F^{\star}}^{2}]\geq\mathrm{Tr}(F^{\star}(N\mathcal{I}_{1}+J_{\pi})^{-1})$. In the likelihood case, $\mathcal{I}_{1}=F^{\star}$, so choose $\pi$ with $J_{\pi}\preceq c_{0}F^{\star}$ on $\Theta_{\operatorname{loc}}$, yielding $\mathrm{Tr}(F^{\star}(NF^{\star}+J_{\pi})^{-1})\geq\mathrm{Tr}(I_{d})/(N+c_{0})\geq c\,d/N$ for a universal $c>0$. ∎
Corollary 6.9 (i.i.d. oracle).
If the oracle draws are i.i.d. $\sim P_{\theta^{\star}}$, then Assumption 6.6 holds with equality in (iii) and $\mathcal{I}_{1}(\theta^{\star})=F^{\star}$. The lower bound (39) reduces to the classical Fisher-information scaling $\Omega(d/N)$.
@@ 7.3 High-probability oracle complexity
Run SGD with constant batch size $b$ and constant stepsize $\eta\leq 1/(4M)$:
@@
Step 3 (sum the descent). Summing from $t=0$ to T1T-1 and using $\mathbb{E}[\|\xi_{t}\|_{2}^{2}\mid\mathcal{F}_{t}]\leq b^{-1}\mathrm{Tr}(F^{\star})$:
$\tfrac{3\eta}{4}\!\sum_{t=0}^{T-1}\|\nabla L(\theta_{t})\|_{2}^{2}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star}),\quad\Delta L:=L(\theta_{0})-L(\theta^{\star}).$
Step 4 (apply Freedman’s inequality). The predictable quadratic variation satisfies $\langle M\rangle_{T}\leq(\eta^{2}/b)\mathrm{Tr}(F^{\star})\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$. The increments $\Delta M_{t+1}=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ have bounded conditional variance $\mathbb{E}[(\Delta M_{t+1})^{2}\mid\mathcal{F}_{t}]\leq\eta^{2}\|\nabla L(\theta_{t})\|_{2}^{2}\cdot b^{-1}\mathrm{Tr}(F^{\star})$, and satisfy $|\Delta M_{t+1}|\leq R$ almost surely for $R:=\eta\,M\,\mathrm{diam}(\mathcal{U})\cdot b^{-1/2}\|F^{\star}\|_{\operatorname{op}}^{1/2}$ (using Cauchy–Schwarz and the bounded-gradient condition on $\mathcal{U}$). By Freedman’s inequality (Lemma A.7), with probability at least $1-\delta$, $M_{T}\leq c\sqrt{\langle M\rangle_{T}\log(1/\delta)}+c^{\prime}R\log(1/\delta)$.
Step 5 (solve for gradient bound). Let $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$. Substituting into Step 3 gives a quadratic inequality in $\sqrt{S_{T}}$; solving:
@@
Classical oracle-complexity results for first-order methods express the iteration count through the Euclidean (Hessian) condition number $\kappa_{H}:=\lambda_{\max}(H^{\star})/\lambda_{\min}(H^{\star})$ and the ambient dimension $d$. In (44), these quantities are replaced by their statistical counterparts: the Fisher condition number $\kappa_{F}=\lambda_{\max}(F^{\star})/\lambda_{\min}(F^{\star})$ and the effective dimension $d_{\operatorname{eff}}(F^{\star})$. The substitution is not cosmetic. A problem may be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and Fisher information share the same eigenstructure, so that difficult curvature directions coincide with informative ones. Conversely, isotropic curvature ($\kappa_{H}=1$) does not guarantee favorable statistical complexity if information is anisotropic. The replacement of $(\kappa_{H},d)$ by $(\kappa_{F},d_{\operatorname{eff}})$ reflects the fact that, once noise geometry is identified, the fundamental bottleneck in stochastic optimization is statistical resolution, not algebraic curvature.
@@
The oracle complexity (44) is governed by two geometry-dependent quantities:
  • $\kappa_{F}=\lambda_{\max}(F^{\star})/\lambda_{\min}(F^{\star})$ is the Fisher condition number—the ratio of maximal to minimal statistical information per coordinate direction. Large $\kappa_{F}$ signals that some parameter directions are much harder to resolve than others.
  • $d_{\operatorname{eff}}(F^{\star})=\mathrm{Tr}(F^{\star})/\lambda_{\max}(F^{\star})$ is the effective dimension (stable rank of $F^{\star}$): the number of statistically distinguishable directions under the curvature-weighted geometry (see Remark 6.2).
The product $\kappa_{F}\,d_{\operatorname{eff}}$ thus measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale. Importantly, (44) is not merely a rate statement: it is a design law that links the practitioner’s sampling budget $N$ to achievable stationarity through intrinsic problem geometry. All modeling freedom has been eliminated—once the loss and sampling design are fixed, the complexity is determined.

### 03 Apr 2026 v2

@@ 6.1 Local Fisher Geometry and a Frozen-Metric Regime
Suppose $F(\theta)$ is Lipschitz on a neighborhood $\mathcal{U}$ of $\theta^{\star}$: $\|F(\theta)-F(\theta^{\star})\|_{\operatorname{op}}\leq L_{F}\|\theta-\theta^{\star}\|_{2}$ for all $\theta\in\mathcal{U}$. If the SGD iterates satisfy $\|\theta_{t}-\theta^{\star}\|_{2}=o(1)$ almost surely, then for every $\epsilon>0$ there exists $t_{0}$ such that for $t\geq t_{0}$, $(1-\epsilon)F^{\star}\preceq F(\theta_{t})\preceq(1+\epsilon)F^{\star}$.
@@
Assumption 6.2 (Local coercivity in the frozen Fisher geometry).
The constant $\mu_{F}$ defined in (35) is strictly positive. This is not automatic from $H^{\star}\succ 0$ and $F^{\star}\succ 0$ alone; it is an explicit compatibility condition between the loss curvature and noise geometry. It holds whenever $H^{\star}$ and $F^{\star}$ are jointly diagonalizable, and more generally whenever $H^{\star}$ and $F^{\star}$ are sufficiently aligned.
@@
The quantity $d_{\operatorname{eff}}(F^{\star})$ is exactly the stable rank of $F^{\star}$ (Tropp, 2015): $d_{\operatorname{eff}}\leq d$ always, with equality iff $F^{\star}$ is isotropic, and $d_{\operatorname{eff}}\to 1$ when a single direction dominates. It replaces $d$ in both the upper bound (Theorem 6.5) and the oracle complexity (Theorem 7.3), so that when Fisher information is concentrated ($d_{\operatorname{eff}}\ll d$), the guarantees are correspondingly tighter than classical Euclidean bounds.
@@ 6.2 Upper Bound
Remark 6.6 (Consistency with the OU viewpoint).
The $1/(Tb)$ scaling in the frozen Fisher norm is consistent with the OU/Lyapunov stationary analysis: both predict the same leading-order dependence on temperature $\tau=\eta/b$ and intrinsic curvature $\mu_{F}$.
@@ 6.3 Lower Bound
The lower bound proceeds in three layers of increasing generality. We first state the cleanest result in the parametric likelihood/Fisher case, then interpret it through oracle information, and finally discuss the extension to general losses.
6.3.1 Layer 1: Parametric Fisher-case lower bound.
Assumption 6.7 (Local parametric regularity for van Trees).
Let $\{P_{\theta}:\theta\in\Theta_{\operatorname{loc}}\}$ be a parametric family on a neighborhood $\Theta_{\operatorname{loc}}$ of $\theta^{\star}$ with densities $p_{\theta}$. Assume:
  1. (i)
    $\theta\mapsto p_{\theta}(z)$ is twice continuously differentiable for a.e. $z$;
  2. (ii)
    the Fisher information $F(\theta):=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}]$ is continuous and positive definite on $\Theta_{\operatorname{loc}}$;
  3. (iii)
    $\int\|s_{\theta}(z)\|^{2}\,p_{\theta}(z)\,dz$ is uniformly bounded on $\Theta_{\operatorname{loc}}$;
  4. (iv)
    there exists a smooth prior density $\pi$ supported on $\Theta_{\operatorname{loc}}$ with finite Fisher information $J_{\pi}:=\int\|\nabla\log\pi(\theta)\|^{2}\pi(\theta)\,d\theta<\infty$.
Theorem 6.8 (Fisher-case lower bound via van Trees).
Suppose Assumption 6.7 holds and the estimator $\hat{\theta}_{N}$ is based on $N$ i.i.d. observations from $P_{\theta^{\star}}$. Then
for a universal constant $c>0$.
Proof.
This is the special case of Proposition 6.10 below with i.i.d. observations (so the oracle information bound holds with equality). We give the self-contained argument in the proof of Proposition 6.10. ∎
6.3.2 Layer 2: Oracle-information interpretation.
The heart of the lower bound is not i.i.d. sampling per se, but bounded predictable information growth. We now restate the result in terms of an abstract oracle transcript.
@@
Under Assumption 6.9, the predictable quadratic variation of the oracle transcript satisfies $\sum_{t=1}^{N}\mathbb{E}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq N\mathcal{I}_{1}(\theta^{\star})$, so the total information available to any $\mathcal{F}_{N}$-measurable estimator is bounded by $N\mathcal{I}_{1}$.
Proposition 6.10 (Lower bound under predictable information budget).
Suppose Assumption 6.7 holds for the parametric family $\{P_{\theta}\}$ and the oracle satisfies Assumption 6.9 with $\mathcal{I}_{1}(\theta^{\star})=F^{\star}$. Additionally, assume that each conditional law of $Z_{t}$ given $\mathcal{F}_{t-1}$ has a density differentiable in $\theta$, so that the conditional Fisher information $I_{t}(\theta\mid\mathcal{F}_{t-1})$ exists and satisfies $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$ a.s. Then every $\mathcal{F}_{N}$-measurable estimator $\hat{\theta}_{N}$ satisfies
@@
The argument extends Theorem 6.8 from i.i.d. observations to the adaptive oracle transcript. The van Trees bound requires an upper bound on the Fisher information $I_{N}(\theta)$ of the joint likelihood of $(Y_{1},\ldots,Y_{N})$ with respect to $\theta$. Three ingredients deliver this bound.
Step 1 (Adaptivity does not inflate per-step information). At step $t$, the query $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, but $Z_{t}\sim P_{\theta}$ is drawn independently of $\mathcal{F}_{t-1}$ given $\theta$. Hence the conditional distribution of $Z_{t}$ given $\mathcal{F}_{t-1}$ depends on $\theta$ only through $\{P_{\theta}\}$, not through the past transcript.
Step 2 (Second-moment bound == per-step Fisher bound). Under correct specification, $Y_{t}=-s_{\theta}(Z_{t})$ is the negative score. The conditional Fisher information of $Z_{t}$ given $\mathcal{F}_{t-1}$ about $\theta$ is $I_{t}(\theta\mid\mathcal{F}_{t-1})=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}\mid\mathcal{F}_{t-1}]=F(\theta)$—exactly the score second-moment matrix. So Assumption 6.9(iii) directly bounds $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$. To extend from $\theta^{\star}$ to all $\theta\in\Theta_{\operatorname{loc}}$, continuity of $F(\theta)$ (Assumption 6.7(ii)) gives $F(\theta)\preceq(1+\epsilon)F^{\star}$ on a sufficiently small $\Theta_{\operatorname{loc}}$; the factor $(1+\epsilon)$ is absorbed into the universal constant $c$.
Step 3 (Chain rule and conclusion). The chain rule for conditional Fisher information applied to the sequential factorization of the joint likelihood gives
$I_{N}(\theta)=\sum_{t=1}^{N}\mathbb{E}_{\theta}\!\bigl[I_{t}(\theta\mid\mathcal{F}_{t-1})\bigr]\preceq N\,(1{+}\epsilon)\,F^{\star}.$
Substituting into van Trees with prior $\pi$ and sending $\epsilon\to 0$ yields (41). ∎
Remark 6.11 (Beyond i.i.d. oracles).
The classical i.i.d. setting satisfies Assumption 6.9 with equality in (iii). The assumption also accommodates exchangeable or conditionally i.i.d. sampling schemes (e.g., de Finetti mixtures) provided the per-step conditional information is bounded by $\mathcal{I}_{1}$. What is required is a predictable information bound on the oracle transcript, not independence across calls. The critical logical step is that the van Trees inequality applies to the joint transcript likelihood, whose total Fisher information is controlled by summing the per-step conditional bounds.
Remark 6.12 (What enters the lower bound).
The lower bound (41) involves $d$ (ambient dimension) and $N$ (oracle budget). The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ enter the upper bound and oracle complexity (Theorems 6.5 and 7.3) but not the lower bound, which is stated in terms of $d$ to give the sharpest minimax comparison. When $d_{\mathrm{eff}}\ll d$, the gap between upper and lower bounds tightens correspondingly.
6.3.3 Layer 3: General-loss / Godambe extension.
Remark 6.13 (Extension to general losses).
In the general-loss case, the same lower-bound form $\Omega(d/N)$ is recovered under local asymptotic conditions that identify the relevant sandwich geometry. Specifically, if the loss satisfies a local asymptotic normality (LAN) condition at $\theta^{\star}$ with information matrix $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$, then the van Trees argument carries through with $F^{\star}$ replaced by $G^{\star}(\theta^{\star})$ and the dual norm adjusted accordingly. The LAN condition is standard for smooth M-estimators under regularity conditions comparable to those already imposed in Section 3; see Le Cam (1986) and van der Vaart (1998). We state this extension as a remark rather than a theorem because the full verification of LAN for the adaptive oracle transcript of Assumption 6.9 requires additional technical conditions (uniform local asymptotic equicontinuity) that we do not develop here.
@@ 7.3 High-probability oracle complexity
Additionally, assume conditional sub-Gaussianity of the noise: for all $v\in\mathbb{R}^{d}$ with $\|v\|=1$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}[\exp(\lambda\,v^{\top}\xi_{t})\mid\mathcal{F}_{t}]\leq\exp(\lambda^{2}\|F^{\star}\|_{\operatorname{op}}/(2b))$ a.s. (This is satisfied, for example, when per-sample gradients are almost surely bounded on $\mathcal{U}$.) Run SGD with constant batch size $b$ and constant stepsize $\eta\leq 1/(4M)$:
@@
Step 3 (sum the descent). Summing from $t=0$ to T1T-1:
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}}{2}\,Q_{T},\quad\Delta L:=L(\theta_{0})-L(\theta^{\star}),$
where $Q_{T}:=\sum_{t<T}\|\xi_{t}\|_{2}^{2}$. To replace the random $Q_{T}$ by a deterministic bound we need concentration, not just conditional moments. Under conditional sub-Gaussianity, $\|\xi_{t}\|_{2}^{2}$ is conditionally sub-exponential with $\mathbb{E}[\|\xi_{t}\|_{2}^{2}\mid\mathcal{F}_{t}]\leq b^{-1}\mathrm{Tr}(F^{\star})$; Bernstein’s inequality for sums of conditionally sub-exponential random variables (see, e.g., Wainwright 2019, Cor. 2.21) gives, with probability at least $1-\delta$, $Q_{T}\leq(T/b)\mathrm{Tr}(F^{\star})+c\sqrt{T\,d}\,(\|F^{\star}\|_{\operatorname{op}}/b)\sqrt{\log(1/\delta)}.$ Since $M\eta^{2}/2\leq\eta/8$ (as $\eta\leq 1/(4M)$), the deviation is $O(\eta\sqrt{T/b}\,\sqrt{d}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)}\,/b)$, which is lower-order relative to the bias term $\Delta L/(\eta T)$ in the final bound. Absorbing it into the constant $C$ (and taking a union bound over the two $\delta$-events, i.e. replacing $\delta$ by $\delta/2$ throughout):
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})\log\!\tfrac{2}{\delta},$
with probability at least $1-\delta$.
Step 4 (sub-Gaussian martingale concentration). Each increment satisfies $\Delta M_{t+1}=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ where $\nabla L(\theta_{t})$ is $\mathcal{F}_{t}$-measurable. By the conditional sub-Gaussianity assumption, the scalar $\langle v,\xi_{t}\rangle$ with $v:=\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|_{2}$ has sub-Gaussian parameter $\sigma_{t}^{2}\leq\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $c_{t}^{2}:=\eta^{2}\|\nabla L(\theta_{t})\|_{2}^{2}\|F^{\star}\|_{\operatorname{op}}/b$. By the standard exponential supermartingale bound for sums of conditionally sub-Gaussian increments (see, e.g., Wainwright 2019, Theorem 2.19), with probability at least $1-\delta$, $M_{T}\leq\sqrt{2\sum_{t<T}c_{t}^{2}\,\log(1/\delta)}=\eta\sqrt{2(\|F^{\star}\|_{\operatorname{op}}/b)\,S_{T}\,\log(1/\delta)}$, where $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$.
Step 5 (solve for gradient bound). Substituting the Step 4 tail bound into Step 3 gives a quadratic inequality in $\sqrt{S_{T}}$; solving:
@@
Classical oracle-complexity results use the Hessian condition number $\kappa_{H}$ and ambient dimension $d$; here they are replaced by the Fisher counterparts $\kappa_{F}$ and $d_{\operatorname{eff}}$. The substitution is not cosmetic: a problem may be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and information share the same eigenstructure; conversely, isotropic curvature does not guarantee favorable statistical complexity.
@@
The product $\kappa_{F}\,d_{\operatorname{eff}}$ in (46) measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale (see Remark 6.3 for the role of $d_{\operatorname{eff}}$ as the stable rank of $F^{\star}$). Equation (46) links the sampling budget NN to achievable stationarity through intrinsic problem geometry.